• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 16:00:26 on November 18, 2000

    Entry number 52326

    keyword=Beam Time Accounting
    Hall A

    11/18/2000 0800-1600
    Experiment #: e99007
    Shift Leader: Huber
    Shift Workers: Z. Chai, Punjabi/Chang

    08000900100011001200130014001500Total Time
    Acceptable Beam
    in Use
    47 51 43 27 11 48 43 2 4.53333333333333
    Beam Available
    but Not in Use
    0 0 0 0 39 0 13 58 1.83333333333333
    Beam Not Available
    or Unacceptable
    13 9 17 33 10 12 4 0 1.63333333333333
    Accelerator Con-
    figuration Chage
    0 0 0 0 0 0 0 0 0
    Experiment
    Ready
    1 1 1 1 15 1 47 5 6.11666666666667
    Planned Config-
    uration Change
    0 0 0 0 0 0 0 0 0
    Unplanned
    Experiment Down
    0 0 0 0 45 0 13 551.88333333333333
    Hall not
    ready reasons
    Target 0 0 0 0 0 0 0 0 0
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 0 0 0 0 0 0 0 0
    Data Ac 0 0 0 0 0 0 0 0 0
    Other 0 0 0 0 0 0 0 0 0

    Comments: Beam increased from 45-50uA to 70-75 uA starting at 9:06. Beam current reduced to 60-65uA near noon, due to high trip rate. 14:45-16:00 - Q3 and then Q1 quenched due to ESR problems with Halls C and then B.

    Data Entry by: Huber, G.